OpenStudy (anonymous):

What is the equation of the line that passes through the points (-2, 3) and (2, 7)?

OpenStudy (anonymous):

x - y = -1 x - y = -2 x - y = - 5 x - y = - 6

OpenStudy (nenadmatematika):

y-3=1(x+2) so x-y=-5

OpenStudy (anonymous):

m = (7-3)/ (2 + 2) = 4/4 = 1

OpenStudy (anonymous):

y - 7 = x - 2 y = x + 5

OpenStudy (anonymous):

First get your slope. (y2-y1)/(x2-x1)

OpenStudy (anonymous):

(y-y1)=m(x-x1). Get your slope (m) And use one of your points for x1 and y1. Then plug all these in to the formula.